OpenStudy (anonymous):

-5 + 14b <= -4 + 15b

OpenStudy (vishweshshrimali5):

b>= -1

OpenStudy (vishweshshrimali5):

b>= -1

OpenStudy (anonymous):

-5+14b<=-4+15b -5+4<=15b-14b -1<=1b -1/1<=b -1<=b or b>=-1